We measured beat noise. With new suspension, the signal is getting worse from ,probably, seismic .
The peak at 3.73 Hz is suppressed, but other mechanical peaks appear.
RCAV is aligned, the efficiency is ~98%
The plot below shows the beat and seismic measurements (x, y, z directions) in arbitrary unit.
The seismometer is placed on the optical table
These peaks have not been seen before, so it seems that our new suspension is not working well, especially
on vertical (z) direction.
We will investigate what cause those peaks, and see if we can fix them.
Front and back Periscope for RCAV are back on the table. The input beam is aligned.
The transmitted beam is coarsely aligned.
Note that the coupling efficiency for RCAV is down to ~ 85%, probably from the change of cavity's position.
I used the periscope to adjust only beam's position and angle. I haven't tried to reposition the lenses yet.
The input beam has multiple reflections from the periscope's top mirror, but beam is quite center on the mirror, not sure what causes this.
Instead of a single beam, I get one bright center beam and top and bottom faint beams ??
The cavities' insulation is back on both cavity. However, the outer box front panel is left open for the vacuum pumping.
We should be able to measure the beat by tomorrow.
cavity is back in chamber and started pumping. Had some trouble with one of the windows. More details and photos later...